# Preserving Privacy in Large Language Models: A Survey on Current Threats and Solutions

### v1

@@ Abstract
Large Language Models (LLMs) represent a significant advancement in artificial intelligence, finding applications across various domains. However, their reliance on massive internet-sourced datasets for training brings notable privacy issues, which are exacerbated in critical domains (e.g., healthcare). Moreover, certain application-specific scenarios may require fine-tuning these models on private data. This survey critically examines the privacy threats associated with LLMs, emphasizing the potential for these models to memorize and inadvertently reveal sensitive information. We explore current threats by reviewing privacy attacks on LLMs and propose comprehensive solutions for integrating privacy mechanisms throughout the entire learning pipeline. These solutions range from anonymizing training datasets to implementing differential privacy during training or inference and machine unlearning after training. Our comprehensive review of existing literature highlights ongoing challenges, available tools, and future directions for preserving privacy in LLMs. This work aims to guide the development of more secure and trustworthy AI systems by providing a thorough understanding of privacy preservation methods and their effectiveness in mitigating risks.
@@ 1 Introduction
Large Language Models (LLMs) represent a significant advancement in artificial intelligence (AI), demonstrating remarkable capabilities in understanding and generating human-like text, and fueling innovation across various domains (Brown et al., 2020a; Touvron et al., 2023; Zhao et al., 2023). However, their reliance on massive internet-sourced datasets for training raises significant privacy concerns considering that these models are prone to memorize segments of their training data (Feldman, 2020; Feldman & Zhang, 2020). This is further exacerbated if we consider certain application-specific scenarios (e.g, healthcare) that may requires the fine-tuning of publicly available models on user private data (Tramèr et al., 2022). Since language models are specifically trained to generate text, this poses a fundamental threat considering that they might inadvertently disclose memorized private data at any moment (Carlini et al., 2019; 2021).
@@
Privacy is a fundamental human right that encompasses the protection of personal information from unauthorized access and misuse (United Nations, 1948; Nissenbaum, 2010). Preserving privacy in the context language language models thus means implementing measures that ensure the protection of personal and sensitive information from being exposed or misused, following the permissions granted by users to use those data, whether implicit or explicit (Brown et al., 2022a). Preserving privacy is not only crucial to preserve users and public trust in AI technologies, but also a legal requirement in many countries such as the EU with the General Data Protection Regulation (GDPR), the US with the Privacy Act, and China with the Personal Information Protection Law and Data Security Law (Weidinger et al., 2021; Solaiman et al., 2019; Veale et al., 2018). These regulations may include additional guarantees, such as the "right to be forgotten", which allows individuals to withdraw their consent and request the deletion of their personal data, posing an additional challenge for LLMs as erasing a data point from a trained model might require retraining the model from scratch or discarding the entire model (Zhang et al., 2024).
@@
Understanding the distinction between privacy and copyright is crucial in this context. Privacy focuses on the protection and control of personal information, ensuring that individuals’ sensitive data remains confidential and is not shared or exploited without their consent. Copyright, on the other hand, pertains to the legal rights of creators over their original works, such as text, music, or art, and governs how these works can be used, distributed, and reproduced by others (Torremans, 2007; Karamolegkou et al., 2023). While privacy is about protecting personal data from unauthorized access, copyright deals with protecting intellectual property and ensuring that right holders receive recognition and compensation for their work. Our focus in this study is on addressing privacy issues related to the potential exposure of personally identifiable information (PII) in LLMs, rather than exploring copyright infringement or intellectual property rights.
@@
This survey aims to provide an holistic view on the problem of preserving privacy in LLMs by combining both the current privacy threats coupled with the current available solutions and tools to implement them. We believe this approach, which presents both aspects in tandem, to be an essential resource for researchers and practitioners seeking a comprehensive understanding of this field. Specifically this survey targets: (i) readers interested in understanding the current privacy risks associated with LLMs (§3); (ii) readers seeking to learn about the available solutions to defend against privacy attacks (§4 - 5); and (iii) readers looking for tools to make their LLMs resistant to privacy attacks (§6).
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners, align=left, minimum width=5em, edge+=darkgray, line width=1pt, s sep=7pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em, font=, where level=2text width=13em, font=, where level=3text width=11em, font=, where level=4text width=10em, font=, [ Preserving Privacy in Large Language Models (LLMs), fill=hidden-gray!70, ver [ Privacy Attacks §3, fill=hidden-red!30 [ Training Data Extraction §3.1, text width=15em, fill=hidden-red!30 [ Non-adversarial extraction §3.1.1, text width=16em, fill=hidden-red!30 [ Carlini et al. (2021; 2023a); Huang et al. (2022); Nasr et al. (2023); Wang et al. (2024), text width=40em, fill=hidden-red!30 ] ] [ Adversarial prompting §3.1.2, text width=16em, fill=hidden-red!30 [ Nasr et al. (2023); Shen et al. (2023); Li et al. (2023b), text width=25em, fill=hidden-red!30 ] ] ] [ Membership Inference §3.2, text width=15em, fill=hidden-red!30 [ Shadow Models §3.2.1, fill=hidden-red!30 [ Shokri et al. (2016); Song & Shmatikov (2019); Hisamoto et al. (2020), text width=32em, fill=hidden-red!30 ] ] [ Thresholds §3.2.2, fill=hidden-red!30 [ Shi et al. (2023); Mireshghallah et al. (2022); Mattern et al. (2023), text width=32em, fill=hidden-red!30 ] ] ] [ Model Inversion §3.3, text width=15em, fill=hidden-red!30 [ Model Output Inversion §3.3.1, text width=15em, fill=hidden-red!30 [ Fredrikson et al. (2015); Pan et al. (2020); Li et al. (2023c); Morris et al. (2023), text width=36em, fill=hidden-red!30 ] ] [ Gradient Inversion §3.3.2, text width=15em, fill=hidden-red!30 [ Zhu et al. (2019); Deng et al. (2021); Balunovic et al. (2022); Gupta et al. (2022), text width=36em, fill=hidden-red!30 ] ] ] ] [ Solutions §4-5, fill=hidden-blue!70 [ Privacy in Data §4, text width=10em, fill=hidden-blue!70 [ Classical Anonymization §4.1.1, text width=15em, fill=hidden-blue!70 [ Zhou et al. (2022); Romanov & Fedotova (2020); Mosallanezhad et al. (2019), text width=34em, fill=hidden-blue!70 ] ] [ Anonymization with DP §5, text width=15em, fill=hidden-blue!70 [ Chatzikokolakis et al. (2013); Krishna et al. (2021); Tong et al. (2024), text width=34em, fill=hidden-blue!70 ] ] ] [ Privacy in Model §5, text width=10em, fill=hidden-blue!70 [ LLMs Differential
Privacy (DP) §5.1 , text width=8em, fill=hidden-blue!70 [ Pre-training with DP §5.1.1, text width=14em, fill=hidden-blue!70 [ Hoory et al. (2021); Yu et al. (2023); Kerrigan et al. (2020); Ganesh et al. (2023), text width=42em, fill=hidden-blue!70 ] ] [ Fine-tuning with DP §5.1.2, text width=14em, fill=hidden-blue!70 [ Yue et al. (2022); Papernot et al. (2019); Bu et al. (2022b); Mattern et al. (2022a), text width=42em, fill=hidden-blue!70 ] ] [ PEFT with DP §5.1.3, text width=11em, fill=hidden-blue!70 [ Yu et al. (2021a); Li et al. (2021; 2023d), text width=21em, fill=hidden-blue!70 ] ] [ Inference with DP §5.2, text width=11em, fill=hidden-blue!70 [ Wu et al. (2023a); Tang et al. (2024); Duan et al. (2023a); Hong et al. (2023); Mai et al. (2023), text width=52em, fill=hidden-blue!70 ] ] ] [ Federated Learning with DP §5.3, text width=16em, fill=hidden-blue!70 [ Central DP §5.3.1, text width=9em, fill=hidden-blue!70 [ Naseri et al. (2020); McMahan et al. (2018); Xu et al. (2023b); Ramaswamy et al. (2020), text width=46em, fill=hidden-blue!70 ] ] [ Local DP §5.3.2, text width=9em, fill=hidden-blue!70 [ Kasiviswanathan et al. (2008); Shokri & Shmatikov (2015); Bebensee (2019), text width=40em, fill=hidden-blue!70 ] ] ] [ Machine Unlearning §5.4, text width=12em, fill=hidden-blue!70 [ Patil et al. (2023); Wu et al. (2023b); Kassem et al. (2023); Eldan & Russinovich (2023); Bourtoule et al. (2019), text width=50em, fill=hidden-blue!70 ] ] ] ] [ Tools & Frameworks §6, fill=hidden-yellow!70 [ Data Anonymization, text width=10em, fill=hidden-yellow!70 [ ARX, OpenData Anonymizer, Ms Presidio, Spark NLP for Healthcare, text width=33em, fill=hidden-yellow!70 ] ] [ Differential Privacy, text width=10em, fill=hidden-yellow!70 [ private-transformers, dp-transformers, Chorus,OpenDP, IBM DiffPrivLib, dp-few-shot-generation, Opacus , text width=48.5em, fill=hidden-yellow!70 ] ] [ Federated Learning, text width=10em, fill=hidden-yellow!70 [ Flower, pfl, TF Federated, PySyft, FATE, PaddleFL, text width=25em, fill=hidden-yellow!70 ] ] ] ]
@@
Figure 1: Taxonomy of Preserving Privacy in Large Language Models. Red indicates various attack techniques, and blue represents current possible solutions to preserve privacy by acting on the training data or model. Finally, in orange we highlight the currently available tools to preserve privacy. All the papers in this survey are also available at the following GitHub repository.
PII Personal Identifiable Information (PII) refers to any data that can be used to identify a specific individual. This may include direct information such as a person’s name, address, email address, social security number, phone number, genome, fingerprints, and face, as well as any other data that, when combined, could lead to the identification of an individual.
The survey is organized as follows and divided into two main categories: privacy attacks3) and solutions4 - 5). The taxonomy in Figure 1 provides a structured view of our survey on preserving privacy in LLMs. For readers unfamiliar with the terminology, we provide a preliminaries section (§2) that explains the basic concepts related to language models, differential privacy, and federated learning.
@@
Privacy attacks encompass training data extraction (§3.1), membership inference (§3.2), and model inversion (§3.3). Training data extraction involves techniques such as non-adversarial extraction (§3.1.1) and adversarial prompting (§3.1.2), where attackers can potentially extract sensitive information from the training data. Membership inference utilizes shadow models (§3.2.1) or thresholds (§3.2.2) to infer whether specific data points were part of the training set. Model inversion includes model output inversion (§3.3.1) and gradient inversion (§3.3.2) methods that can reconstruct the input data from model outputs.
@@
Solutions to these privacy challenges are categorized into privacy in data (§4) and privacy in model (§5). Privacy in data includes classical anonymization techniques to anonymize data before it is used in training (§4.1.1) and anonymization with differential privacy (DP) (§5), which ensures that individual data points cannot be distinguished from the dataset. Privacy in model encompasses the application of DP during different phases such as pre-training (§5.1), fine-tuning (§5.1.2), Parameter-Efficient Fine-Tuning (PEFT) (§5.1.3), and inference (§5.2). Federated learning with DP (§5.3) is implemented to maintain privacy across distributed datasets, and machine unlearning (§5.4) involves techniques to effectively remove specific data from trained models. Finally, we present available tools and frameworks to implement privacy in language models (§6).
@@
In summary, this survey serves as a critical resource for researchers, and practitioners providing insights into the current privacy threats in LLMs and the currently available solutions to address these challenges. By systematically categorizing privacy issues and their corresponding mitigation strategies, this paper not only highlights the importance of preserving privacy in LLMs but also paves the way for future advancements in the field. All the papers in this survey are also available at the following GitHub repository [^1] https://github.com/michele17284/Awesome-Privacy-Preserving-LLMs.
@@ 2 Preliminaries
In this section, we present preliminary concepts essential for comprehending the subsequent content of the document. Readers already acquainted with these concepts may proceed directly to the next section.
Machine Learning
Machine learning involves the development of algorithms that can learn from and make predictions or decisions based on data by recognizing patterns in it, utilizing various techniques that include both linear and nonlinear models. Classical machine learning often relies on handcrafted features that require domain knowledge to be created. Classical machine learning models include decision trees, support vector machines, linear regression, and ensemble methods like random forests, designed for specific data types and tasks. Models are composed of parameters, which are what is learned in the learning process. These are variables that the learning algorithm adjusts during training to fit the model to the data as closely as possible. Parameters can be weights, as they affect how much of the input is passed between neurons (as a multiplication factor), or biases, shifting the outputs along the value axis (as an addition factor). Network parameters are very influential and their magnitude defines the model complexity and its capacity to recognize more intricate patterns in the data and generalize better. More complex networks fall under deep learning, see 2. There are three different types of learning: supervised learning, where there is a ground truth, and the model tries to output an answer as close as possible to that; unsupervised learning, where there is no answer and the model groups inputs based on some similarity; reinforcement learning, where the model is educated through punishments and rewards. The most common form of learning is supervised learning, which is the most present in this survey. Learning parameters is oftentimes done using some form of gradient descent. The gradient descent algorithm starts at a random point (random initialization of parameters), computes the gradient of the optimized function (loss), and then updates the parameters to reduce the gradient. This is iteratively repeated until a local minimum is reached. This process is composed of three phases: i) feed-forward, the input is fed to the model to produce the output; ii) an error is computed by comparing the actual output with the expected output; iii) the gradient w.r.t. this error is used to update the parameters.
Deep Learning
Deep learning, differently from machine learning, leverages complex neural networks to learn directly from vast volumes of data, without the need for handcrafted features. This approach allows for the automated extraction of intricate patterns and the learning of non-linear relationships within the data. Deep learning architectures are all variations of neural networks. The name comes from the initial idea of mimicking a human brain. They are often characterized by their depth, consisting of multiple layers of neurons, each contributing to the model’s ability to perform tasks such as image and speech recognition, language modeling, and more, with unprecedented accuracy. Each layer contains a set of parameters or weights that adjust the strength (the "weight" precisely) of the signal being transmitted between neurons by multiplying it. Most deep learning networks are trained with gradient descent to adjust network parameters, effectively minimizing the difference between actual and predicted outcomes through an iterative process. This includes both feed-forward mechanisms, where data flows through the network to produce an output, and backpropagation, where errors are propagated back to network parameters to update the model. Usually the gradients are not averaged over the whole dataset (Batch Gradient Descent), because especially on large datasets (which is often the case with DNNs) it is not very efficient. More often, Stochastic Gradient Descent is used, which computes the gradients over a mini-batch of the actual dataset (Zhang, 2004; Ruder, 2017). Deep learning relies on large datasets and much computational power but significantly outperforms classical models in complex tasks.
Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) represent a significant advancement in artificial intelligence, finding applications across various domains. However, their reliance on massive internet-sourced datasets for training brings notable privacy issues exacerbated in critical domains (e.g., healthcare). Moreover, certain application-specific scenarios may require fine-tuning these models on private data. This survey critically examines the privacy threats associated with LLMs, emphasizing the potential for these models to memorize and inadvertently reveal sensitive information. We explore current threats by reviewing privacy attacks on LLMs and propose comprehensive solutions for integrating privacy mechanisms throughout the entire learning pipeline. These solutions range from anonymizing training datasets to implementing differential privacy during training or inference and machine unlearning after training. Our comprehensive review of existing literature highlights ongoing challenges, available tools, and future directions for preserving privacy in LLMs. This work aims to guide the development of more secure and trustworthy AI systems by providing a thorough understanding of privacy preservation methods and their effectiveness in mitigating risks.
@@ 1 Introduction
Large Language Models (LLMs) represent a significant advancement in artificial intelligence (AI), demonstrating remarkable capabilities in understanding and generating human-like text, and fueling innovation across various domains (Brown et al., 2020; Touvron et al., 2023; Zhao et al., 2023). However, their reliance on massive internet-sourced datasets for training raises significant privacy concerns, considering that these models are prone to memorize segments of their training data (Feldman, 2020; Feldman & Zhang, 2020). This is further exacerbated if we consider certain application-specific scenarios (e.g. healthcare) that may require fine-tuning publicly available models on user private data (Tramèr et al., 2024a). Since language models are specifically trained to generate text, this poses a fundamental threat considering that they might inadvertently disclose memorized private data at any moment (Carlini et al., 2019; 2021).
@@
Privacy is a fundamental human right that encompasses the protection of personal information from unauthorized access and misuse (United Nations, 1948; Nissenbaum, 2010). Preserving privacy in the context of language models thus means implementing measures that ensure the protection of personal and sensitive information from being exposed or misused, following the permissions granted by users to use those data, whether implicit or explicit (Brown et al., 2022a). Preserving privacy is not only crucial to preserve users and public trust in AI technologies but also a legal requirement in many countries such as the EU with the General Data Protection Regulation (GDPR), the US with the Privacy Act, and China with the Personal Information Protection Law and Data Security Law (Weidinger et al., 2021; Solaiman et al., 2019; Veale et al., 2018). These regulations may include additional guarantees, such as the "right to be forgotten", which allows individuals to withdraw their consent and request the deletion of their personal data, posing an additional challenge for LLMs as erasing a data point from a trained model might require retraining the model from scratch or discarding the entire model (Zhang et al., 2024b).
@@
Understanding the distinction between privacy and copyright is crucial in this context. Privacy focuses on protecting and controlling personal information, ensuring that individuals’ sensitive data remains confidential and is not shared or exploited without their consent. Copyright, on the other hand, pertains to the legal rights of creators over their original works, such as text, music, or art, and governs how these works can be used, distributed, and reproduced by others (Torremans, 2007; Karamolegkou et al., 2023). While privacy is about protecting personal data from unauthorized access, copyright protects intellectual property and ensures rights holders receive recognition and compensation for their work. This study focuses on addressing privacy issues related to the potential exposure of personally identifiable information (PII) in LLMs rather than exploring copyright infringement or intellectual property rights.
PII Personal Identifiable Information (PII) refers to any data that can be used to identify a specific individual. This may include direct information such as a person’s name, address, email address, social security number, phone number, genome, fingerprints, and face, as well as any other data that, when combined, could lead to the identification of an individual (National Institute of Standards and Technology, 2017; General Data Protection Regulation, 2016).
@@
This survey aims to provide a holistic view of the problem of preserving privacy in LLMs by combining the current privacy threats and the current available solutions and tools to implement them. We believe this approach presents both aspects in tandem and is an essential resource for researchers and practitioners seeking a comprehensive understanding of this field. Specifically, this survey targets (i) readers interested in understanding the current privacy risks associated with LLMs (§3); (ii) readers seeking to learn about the available solutions to defend against privacy attacks (§4 - 5); and (iii) readers looking for tools to make their LLMs resistant to privacy attacks (§6).
The survey is organized as follows and divided into two main categories: privacy attacks3) and solutions4 - 5). The taxonomy in Figure 1 provides a structured view of our survey on preserving privacy in LLMs. We provide a preliminaries section for readers unfamiliar with the terminology (§2) that explains the basic concepts related to language models, differential privacy, and federated learning.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners, align=left, minimum width=5em, edge+=darkgray, line width=1pt, s sep=7pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em, font=, where level=2text width=13em, font=, where level=3text width=11em, font=, where level=4text width=10em, font=, [ Preserving Privacy in Large Language Models (LLMs), fill=hidden-gray!70, ver [ Privacy Attacks §3,text width=9em, fill=hidden-red!30 [ Training Data Extraction §3.1, text width=13.5em, fill=hidden-red!30 [ Non-adversarial extraction §3.1.1, text width=15em, fill=hidden-red!30 ] [ Adversarial prompting §3.1.2, text width=13.5em, fill=hidden-red!30 ] ] [ Membership Inference §3.2, text width=13em, fill=hidden-red!30 [ Thresholds §3.2.1,text width=8em, fill=hidden-red!30 ] ] [ Model Inversion and Stealing §3.3, text width=15em, fill=hidden-red!30 [ Model Output Inversion §3.3.1, text width=14em, fill=hidden-red!30 ] [ Gradient Inversion §3.3.2, text width=11.5em, fill=hidden-red!30 ] [ Model Stealing §3.3.3, text width=10em, fill=hidden-red!30 ] ] [ Privacy Threats at Inference Time §3.4, text width=17.5em, fill=hidden-red!30 ] ] [ Solutions §4-5,text width=7em, fill=hidden-blue!70 [ Privacy in Data §4, text width=9em, fill=hidden-blue!70 [ Classical Anonymization §4.1.1, text width=14em, fill=hidden-blue!70 ] [ Anonymization with DP §5, text width=14em, fill=hidden-blue!70 ] ] [ Privacy in Model §5, text width=9em, fill=hidden-blue!70 [ LLMs Differential
Privacy (DP) §5.1 , text width=8em, fill=hidden-blue!70 [ Pre-training with DP §5.1.1, text width=12.5em, fill=hidden-blue!70 ] [ Fine-tuning with DP §5.1.2, text width=12.5em, fill=hidden-blue!70 ] [ PEFT with DP §5.1.3, text width=10em, fill=hidden-blue!70 ] [ Inference with DP §5.2, text width=10.5em, fill=hidden-blue!70 ] ] [ Federated Learning with DP §5.3, text width=15em, fill=hidden-blue!70 ] [ Machine Unlearning §5.4, text width=11.5em, fill=hidden-blue!70 ] [ Homomorphic Encryption §5.5, text width=14em, fill=hidden-blue!70 ] ] ] [ Tools & Frameworks §6,text width=11em, fill=hidden-yellow!70 [ Data Anonymization, text width=9.5em, fill=hidden-yellow!70 [ ARX, OpenData Anonymizer, Presidio,
Spark NLP for Healthcare, text width=19em, fill=hidden-yellow!70 ] ] [ Differential Privacy, text width=8.5em, fill=hidden-yellow!70 [ private-transformers, dp-transformers, Chorus,OpenDP,
DiffPrivLib,dp-few-shot-generation, Opacus, text width=24.5em, fill=hidden-yellow!70 ] ] [ Federated Learning, text width=8.5em, fill=hidden-yellow!70 [ Flower, pfl, TF Federated, PySyft, FATE, PaddleFL, text width=23em, fill=hidden-yellow!70 ] ] ] ]
@@
Figure 1: Taxonomy of Preserving Privacy in Large Language Models. Red indicates various attack techniques, and blue represents current possible solutions to preserve privacy by acting on the training data or model. Finally, in orange we highlight the currently available tools to preserve privacy.
@@
Privacy attacks discuss how it’s possible to extract training data from a model and encompass training data extraction (§3.1), membership inference (§3.2), and model inversion and stealing (§3.3). Training data extraction involves techniques such as non-adversarial extraction (§3.1.1) and adversarial prompting (§3.1.2), where attackers can potentially extract sensitive information from the training data. Membership inference (§3.2) utilizes shadow models or thresholds (§3.2.1) to infer whether specific data points were part of the training set. Model inversion includes model output inversion (§3.3.1) and gradient inversion (§3.3.2) methods that can reconstruct the input data from model outputs while Model stealing is proposed to reconstruct a model’s parameters from its output (§3.3.3). However, as prompting LLMs with private input – such as in-context demonstrations – is an increasingly popular approach, threats at inference time are also discussed 3.4.
@@
Solutions to these privacy challenges are categorized into privacy in data (§4) and privacy in model (§5). Privacy in data includes classical anonymization techniques to anonymize data before it is used in training (§4.1.1) and anonymization with differential privacy (DP) (§5), which ensures that individual data points cannot be distinguished from the dataset. Privacy in model encompasses the application of DP during different phases such as pre-training (§5.1), fine-tuning (§5.1.2), Parameter-Efficient Fine-Tuning (PEFT) (§5.1.3), and inference (§5.2). Federated learning with DP (§5.3) is implemented to maintain privacy across distributed datasets, and machine unlearning (§5.4) involves techniques to effectively remove specific data from trained models. Homomorphic Encryption (§5.5) involves cryptographic methods that allow computation on encrypted data without decrypting it. Finally, we present available tools and frameworks to implement privacy in language models (§6).
@@
In summary, this survey serves as a critical resource for researchers and practitioners, providing insights into the current privacy threats in LLMs and the currently available solutions to address these challenges. By systematically categorizing privacy issues and their corresponding mitigation strategies, this paper highlights the importance of preserving privacy in LLMs and paves the way for future advancements in the field. All the papers in this survey are also available at the following GitHub repository [^1] https://github.com/michele17284/Awesome-Privacy-Preserving-LLMs.
@@ 2 Preliminaries
In this section, we present preliminary concepts essential for comprehending the subsequent content of the document. Readers already acquainted with these concepts may proceed directly to the next section. For a comprehensive introduction to the concepts of Machine Learning and Deep Learning, we direct the reader to Bishop (2006) and Goodfellow et al. (2016).
In this survey, we use the term large language model (LLM) to indicate any model trained on broad textual data that can be adapted (e.g., via fine-tuning or in-context learning) to a wide range of downstream tasks (Bommasani et al., 2021). While the focus of the survey is mostly on generative LLMs (Zhao et al., 2023)— models that are capable of solving a broad set of tasks by generating text, such as GPT-3 and GPT-4— it is not uncommon to encounter smaller models like BERT in our discussion. We have chosen to include some of these models when we believe that their associated findings, both in terms of potential threats and proposed solutions, remain still relevant and generalizable to larger models nowadays.
2.1 Large Language Models